Suppose the short run production function can be represented by

Q = 60,000L2– 1000L3. Then, determine

a. The level labor employment that maximizes the level of output

b. The level of employment that maximizes APL and the maximum APL

1
Expert's answer
2021-12-17T08:40:36-0500

a. The level labor employment that maximizes the level of output

The total output will be maximized when the marginal product is equal to zero.

"MPL=\dfrac{dQ}{dL}\\[0.3cm]
MPL = 120,000L– 3000L^2=0\\[0.3cm]
L(120,000– 3000L)=0\\[0.3cm]
120,000– 3000L=0\\[0.3cm]
3000L=120,0000\\[0.3cm]
L^*=\dfrac{120,000}{3000}=\boxed{40}"

b. The level of employment that maximizes APL and the maximum APL

The average product is maximized when the marginal product is equal to the average product.


The average product is

"AP=\dfrac{Q}{L}\\[0.3cm]
AP=60,000L-1000L^2"

Therefore

"60,000L-1000L^2=120,000L-3000L^2\\[0.3cm]
4000L^2=60000L\\[0.3cm]
4000L=60000\\[0.3cm]
L=\dfrac{60000}{4000}=\boxed{15}"

The maximum average product is equal to

"AP=60,000(15)-1000(15)^2\\[0.3cm]
AP=\boxed{675,000}"